MohamedKhan:

Let point C be between V and W on 2004-01-01-02-00_files/i0350000.jpg. Given that VW = 61, VC = z + 13, and CW = z + 8, solve for

jhonyy9:

so given that C is between V and W and that VW =61,VC = z+13 and CW=z+8 solve for ... ? for what ? suppose for z from these given details you can writing VW = VC+CW so 61=z+13+z+8 61=2z+21 61-21=2z 40=2z divide both sides by 2 20=z hope helped and this is understandably easy